# StreamingThinker: Large Language Models Can Think While Reading

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities in chain of thought (CoT) reasoning. However, the current LLM reasoning paradigm initiates thinking only after the entire input is available, which introduces unnecessary latency and weakens attention to earlier information in dynamic scenarios. Inspired by human cognition of thinking while reading, we first design a streaming thinking paradigm for LLMs, where reasoning unfolds in the order of input and further adjusts its depth once reading is complete. We instantiate this paradigm with StreamingThinker, a framework that enables LLMs to think while reading through the integration of streaming CoT generation, streaming-constraint training, and streaming parallel inference. Specifically, StreamingThinker employs streaming reasoning units with quality control for CoT generation, enforces order-preserving reasoning through streaming attention masks and position encoding, and leverages parallel KV caches that decouple input encoding from reasoning generation, thereby ensuring alignment and enabling true concurrency. We evaluate StreamingThinker on the Qwen3 model family across math reasoning, logical reasoning, and context-based QA reasoning tasks. Experimental results show that the StreamingThinker preserves performance comparable to batch thinking, while yielding an 80% reduction in token waiting before the onset of reasoning and a more than 60% reduction in time-level latency for producing the final answer, demonstrating the effectiveness of the streaming paradigm for LLM reasoning. Code will be released at this repository.
@@ 1 Introduction
Large language models (LLMs) have shown impressive reasoning capabilities, as exemplified by systems like OpenAI-o1 (jaech2024openai) and DeepSeek-R1 (guo2025deepseek). Yet, most current approaches follow a batch thinking paradigm, in which reasoning begins only after the entire input context has been received. This paradigm is problematic in scenarios that demand timely responses or dynamic information processing. First, waiting for the full input before reasoning introduces unnecessary latency. Second, as the input increases, attention to earlier information becomes diluted due to the growing disconnect between reasoning steps and their relevant context (liu2024lost; levy2024same; zhang2025attention). This weakens coherence and raises the risk of hallucination. To compensate, LLMs often rely on longer chains of thought (CoT) (wei2022chain; yeo2025demystifying; chen2025towards) or repeated self-refinement (wangself; ling2023deductive; madaan2023self) to re-focus, which in turn raise computational costs and inflate token usage.
In contrast, human reasoning often unfolds in an immediate and streaming manner. Research in psychology and cognitive science shows that during reading, humans process incoming information instantaneously, including text decoding, meaning construction, background knowledge activation, integrative reasoning, and actively generating inferences to build a coherent understanding (kintsch1988role; graesser1994constructing). This “thinking while reading” mechanism not only enhances processing efficiency, but also allows reasoning to occur closely alongside the relevant context, minimizing cognitive lag and mitigating the risk of coherence degradation.
@@
Accordingly, we propose StreamingThinker, a framework that instantiates the streaming thinking paradigm. The StreamingThinker integrates a streaming CoT generation pipeline with training and inference frameworks that adapt LLMs to the streaming paradigm. The generation pipeline inserts boundary tokens to the inputs to define minimal reasoning units, prompting LLMs to generate serialized reasoning segments for each unit that are reconstructed into incremental content, filtered by quality evaluation, with reasoning depth controlled through token intervention (wu2025effectively). To support streaming training, StreamingThinker introduces two modifications: a streaming attention mask that restricts each reasoning step to past and current input, and a streaming position encoding that independently indexes input and reasoning tokens from zero to eliminate positional contention, ensuring alignment with associated inputs. For inference, StreamingThinker employs parallel KV caches that decouple input encoding from reasoning generation and merge only during cross-attention, enabling true thinking while reading. We conduct a comprehensive evaluation on Qwen3 model family (yang2025qwen3), covering diverse tasks such as math reasoning, logical reasoning, and context-based QA reasoning. Experimental results indicate that StreamingThinker achieves reasoning performance on par with batch thinking, yet reduces token-level waiting before reasoning by 80% and overall answer latency by over 60%.
@@ 2 Streaming Thinking Paradigm
Human streaming cognition involves two complementary processes: rapidly generating and updating representations as input arrives, and subsequently performing global integration to transform local, shallow understanding into holistic, deep comprehension (kintsch1988role).
@@
In the batch thinking paradigm, all input is available simultaneously, so the order of context and question is often overlooked. Yet prior work shows that their relative order can influence reasoning (chen2024premise; wei2024unveiling; xie2024order), an effect amplified in streaming scenarios. In human reading, two natural input orders commonly occur. In the first, the question is presented before the context, enabling the reader to establish targeted associations as subsequent information is processed. In the second, the context precedes the question, in which case the question remains unavailable during streaming reasoning and the reader must rely solely on the context itself to construct plausible inferences. To approximate these scenarios, our streaming thinking paradigm explicitly distinguishes between the two orders and examines their impact on model reasoning. Some examples are provided in Appendix B.
@@ 3.1 Streaming CoT Generation
To further enforce sequential alignment, a higher-parameter teacher model reconstructs the generated reasoning. Once all sentence-level reasoning traces are generated, they are evaluated using granularity score and sequential consistency score. Passing samples are enhanced with token-level intervention to generate depth-controlled reasoning variants. Samples that fail the evaluation are regenerated, and those still failing under Pass@2 (chen2021evaluating) metric are discarded.
@@
We propose two evaluation metrics: the granularity score measures the fine-grained alignment of the streaming CoT, while the sequential consistency score assesses whether the reasoning proceeds in a streaming, order-preserving way. The granularity score is defined as the ratio between the number of boundary tokens in the input and those in the output: , where and denote the counts of boundary tokens in the input and output, respectively. When the granularity score is equal to 1, the reasoning matches the input in boundary count, suggesting ideal alignment. The consistency score is defined as the similarity between the input sentence and reasoning sentences, i.e., , where and denote the embedding vectors of the reasoning sentences and the input sentence , respectively. We use SentenceBERT (reimers2019sentence) for sequential consistency calculation.333In cases where a single input sentence corresponds to multiple reasoning sentences, we regard them collectively as one segment. We provide the similarity map between the input and the reasoning in the Appendix C.5.
@@ 3.2 Streaming Training Framework
The RoPE (su2024roformer) of LLMs represents relative positions by rotating queries and keys, with the attention between a reasoning token and an input token expressed as , where is the input length and are their positional IDs, and is the rotary matrix. However, in streaming scenarios, the concurrent generation of output and reception of input induces positional contention in the encoding space (tong2025llm; guo2024decoder). To address this issue, we assign independent position IDs to input and reasoning tokens, both starting from zero. Formally, the positional IDs of reasoning token and input token are both set to , yielding . This design removes positional contention in streaming parallel processing. Furthermore, identical position IDs ensure that, during streaming reasoning, a reasoning sentence is positioned nearest to its associated input sentence and distant from others, which conforms to the essential principle of streaming alignment.
@@ 4.1 Experimental Settings
To evaluate StreamingThinker, we conduct a comprehensive assessment across three representative reasoning tasks: math reasoning, logical reasoning, and context-based QA reasoning. For math reasoning, we selected GSM-symbolic (mirzadehgsm) and MetamathQA (yumetamath). For logical reasoning, we utilized LogicNLI (tian2021diagnosing) and ProofWriter (tafjord2020proofwriter). Finally, for context-based QA reasoning, PubMedQA (jin2019pubmedqa) and HotpotQA (yang2018hotpotqa) were employed. All datasets were partitioned into dedicated training and testing sets, with detailed specifications provided in the Appendix D.
@@
The evaluation of streaming scenarios can be viewed as a trade-off between performance and latency. For reasoning performance, we adopt Pass@1 score as the accuracy metric to measure the model’s ability to successfully solve problems. Latency is assessed at two levels: token latency and time latency. At the token level, we use token-to-first-token (TTFT) to measure how many input tokens must be observed before the model begins reasoning.444This is similar with time-to-first-token, but measured in terms of token count. For time latency, we set the LLM’s input speed as the average human speaking rate about 150 words per minute (geva2006reading; jacewicz2010between), and define the waiting time until the first answer token as the latency.555Detailed definitions and additional evaluation metrics are provided in Appendix F.
@@ 5 Discussion
Prior studies (fan2025missing; laban2025llms) have highlighted the risks of reasoning over incomplete inputs. This phenomenon reflects the inadequacy of batch-processing LLMs when facing local contexts, as their training objectives align with global visibility rather than incremental inference. However, Streaming Thinker circumvents these pitfalls through three distinct mechanisms. First, global information is deferred rather than lost. Unlike scenarios where critical conditions are permanently removed, our framework merely shifts the timing of acquisition, ensuring the model incorporates the full global context after the streaming phase. Second, we employ a conservative reasoning strategy. Instead of attempting premature complex reflection, the model concentrates on shallow reasoning” (e.g., intermediate calculations and entity tracking) during the streaming phase. This functions as an incremental pre-processing step that simplifies raw context. Third, this behavior is enforced via a specialized streaming training paradigm. Unlike batch models that rigidly apply full-context patterns to partial inputs—often falling into the over-thinking” trap—our model is explicitly adapted to local scopes, learning to process available information without jumping to erroneous conclusions.
@@ 6 Related Work
Prior studies on efficient reasoning in LLMs have mainly focused on three directions: token compression (aggarwal2025l1; xia2025tokenskip; zhang2025lightthinker), structural quantization and pruning (liu2025quantization; srivastava2025towards; zhao2025skipgpt), and efficient decoding (pan2025learning; liao2025reward). Token compression methods such as Tokenskip (xia2025tokenskip) condense CoT into fewer tokens to reduce the inference cost. Structural approaches leverage quantization or pruning to compress model parameters for efficient reasoning such as (srivastava2025towards). Efficient decoding has been explored through parallel sampling of generation paths (pan2025learning) or by using smaller models for speculative prediction (liao2025reward) to reduce inference latency.
In contrast, our work introduces streaming processing as a complementary dimension of efficiency, allowing reasoning to proceed concurrently with input processing to reduce response latency.
@@
Recent research on streaming LLMs has focused on three directions: architecture adaptation (tong2025llm; raffel2024simultaneous; guo2024decoder), latency control (ma2018stacl; ahmed2025non; cheng2025seed), and modality adaptation (chen2024videollm; chen2025livecc; xie2024mini). Architecture adaptation addresses the mismatch between decoder-only transformers and streaming settings (tong2025llm), where issues such as positional interference and redundant re-encoding are mitigated through recurrent states, and modified attention mechanisms. Latency control emphasizes fine-grained alignment between input and output, ranging from fixed policies (e.g., wait-k (ma2018stacl)) to adaptive scheduling that optimize the accuracy–latency balance. Modality adaptation extends streaming LLMs beyond text to speech recognition (guo2024decoder), speech translation (cheng2025seed), and video understanding (chen2024videollm) by integrating modality-specific encoders and synchronization mechanisms. From a reasoning perspective, recent studies (xie2025interleaved; chiang2025stitch; xie2025mini) have explored alternating reasoning and generation to approximate streaming operation. Our work differs by explicitly modeling thinking while reading, enabling reasoning to evolve concurrently with incremental input.
@@ B.1 Cognitive Foundations of Streaming Thinking
Our proposed StreamingThinker architecture draws structural inspiration from established models of human discourse comprehension, specifically the Construction-Integration (CI) model proposed by Kintsch (kintsch1988role). The CI model posits that comprehension occurs in two alternating cycles:
@@
Inspired by this cognitive process, we proposed in the main text a Streaming Thinking Paradigm that enables large language models to reason concurrently with incremental input. Our streaming phase corresponds to the construction phase, where the model performs ”shallow reasoning” (e.g., entity tracking, intermediate calculation) to process incoming chunks and maintain local coherence (graesser1994constructing). Our final reasoning phase corresponds to the integration phase, where the model utilizes the fully accumulated context (KV cache) to synthesize a globally consistent answer. This theoretical alignment explains why our model avoids the ”hallucination” pitfalls of premature guessing: like a human reader, it defers the final integration of complex causal chains until the necessary global information is available.
@@ C.1 Does Incremental Sentence Input Work?
A straightforward idea is to feed sentences into LLMs sequentially and expect sentence-by-sentence reasoning based on the accumulated input. However, our experiments show that when given only a single sentence without streaming fine-tuning, the original reasoning model tends to overthink (sui2025stop). In other words, although this sentence-level data generation method appears to align better with the streaming property, the vanilla model lacks such generative capability and exhibits reduced controllability in CoT generation.
@@ C.4 Quality Evaluation of Streaming CoT
where and denote the embedding vectors of reasoning sentence and input sentence , respectively. We employ SentenceBERT (reimers2019sentence) to obtain embeddings. A higher score reflects that the reasoning faithfully preserves both the order and semantic content of the input stream. In cases where multiple reasoning sentences correspond to one input sentence, we aggregate them as a single segment.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities in chain of thought (CoT) reasoning. However, the current LLM reasoning paradigm initiates thinking only after the entire input is available, which introduces unnecessary latency and weakens attention to earlier information in dynamic scenarios. Inspired by human cognition of thinking while reading, we first design a streaming thinking paradigm for LLMs, where reasoning unfolds in the order of input and further adjusts its depth once reading is complete. We instantiate this paradigm with StreamingThinker, a framework that enables LLMs to think while reading through the integration of streaming CoT generation, streaming-constraint training, and streaming parallel inference. Specifically, StreamingThinker employs streaming reasoning units with quality control for CoT generation, enforces order-preserving reasoning through streaming attention masks and position encoding, and leverages parallel KV caches that decouple input encoding from reasoning generation, thereby ensuring alignment and enabling true concurrency. We evaluate StreamingThinker on the Qwen3 model family across math reasoning, logical reasoning, and context-based QA reasoning tasks. Experimental results show that the StreamingThinker preserves performance comparable to batch thinking, while yielding an 80% reduction in token waiting before the onset of reasoning and a more than 60% reduction in time-level latency for producing the final answer, demonstrating the effectiveness of the streaming paradigm for LLM reasoning. Code is publicly available at this repository.
@@ 1 Introduction
Large language models (LLMs) have shown impressive reasoning capabilities, as exemplified by systems like OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). Yet, most current approaches follow a batch thinking paradigm, in which reasoning begins only after the entire input context has been received. This paradigm is problematic in scenarios that demand timely responses or dynamic information processing. First, waiting for the full input before reasoning introduces unnecessary latency. Second, as the input increases, attention to earlier information becomes diluted due to the growing disconnect between reasoning steps and their relevant context (Liu et al., 2024; Levy et al., 2024; Zhang et al., 2025b). This weakens coherence and raises the risk of hallucination. To compensate, LLMs often rely on longer chains of thought (CoT) (Wei et al., 2022; Yeo et al., 2025; Chen et al., 2025b) or repeated self-refinement (Wang et al., 2023; Ling et al., 2023; Madaan et al., 2023) to re-focus, which in turn raise computational costs and inflate token usage.
In contrast, human reasoning often unfolds in an immediate and streaming manner. Research in psychology and cognitive science shows that during reading, humans process incoming information instantaneously, including text decoding, meaning construction, background knowledge activation, integrative reasoning, and actively generating inferences to build a coherent understanding (Kintsch, 1988; Graesser et al., 1994). This “thinking while reading” mechanism not only enhances processing efficiency, but also allows reasoning to occur closely alongside the relevant context, minimizing cognitive lag and mitigating the risk of coherence degradation.
@@
Accordingly, we propose StreamingThinker, a framework that instantiates the streaming thinking paradigm. The StreamingThinker integrates a streaming CoT generation pipeline with training and inference frameworks that adapt LLMs to the streaming paradigm. The generation pipeline inserts boundary tokens to the inputs to define minimal reasoning units, prompting LLMs to generate serialized reasoning segments for each unit that are reconstructed into incremental content, filtered by quality evaluation, with reasoning depth controlled through token intervention (Wu et al., 2025). To support streaming training, StreamingThinker introduces two modifications: a streaming attention mask that restricts each reasoning step to past and current input, and a streaming position encoding that independently indexes input and reasoning tokens from zero to eliminate positional contention, ensuring alignment with associated inputs. For inference, StreamingThinker employs parallel KV caches that decouple input encoding from reasoning generation and merge only during cross-attention, enabling true thinking while reading. We conduct a comprehensive evaluation on Qwen3 model family (Yang et al., 2025), covering diverse tasks such as math reasoning, logical reasoning, and context-based QA reasoning. Experimental results indicate that StreamingThinker achieves reasoning performance on par with batch thinking, yet reduces token-level waiting before reasoning by 80% and overall answer latency by over 60%.
@@ 2 Streaming Thinking Paradigm
Human streaming cognition involves two complementary processes: rapidly generating and updating representations as input arrives, and subsequently performing global integration to transform local, shallow understanding into holistic, deep comprehension (Kintsch, 1988).
@@
In the batch thinking paradigm, all input is available simultaneously, so the order of context and question is often overlooked. Yet prior work shows that their relative order can influence reasoning (Chen et al., 2024b; Wei et al., 2024; Xie, 2024), an effect amplified in streaming scenarios. In human reading, two natural input orders commonly occur. In the first, the question is presented before the context, enabling the reader to establish targeted associations as subsequent information is processed. In the second, the context precedes the question, in which case the question remains unavailable during streaming reasoning and the reader must rely solely on the context itself to construct plausible inferences. To approximate these scenarios, our streaming thinking paradigm explicitly distinguishes between the two orders and examines their impact on model reasoning. Some examples are provided in Appendix B.
@@ 3.1 Streaming CoT Generation
To further enforce sequential alignment, a higher-parameter teacher model reconstructs the generated reasoning. Once all sentence-level reasoning traces are generated, they are evaluated using granularity score and sequential consistency score. Passing samples are enhanced with token-level intervention to generate depth-controlled reasoning variants. Samples that fail the evaluation are regenerated, and those still failing under Pass@2 (Chen et al., 2021) metric are discarded.
@@
We propose two evaluation metrics: the granularity score measures the fine-grained alignment of the streaming CoT, while the sequential consistency score assesses whether the reasoning proceeds in a streaming, order-preserving way. The granularity score is defined as the ratio between the number of boundary tokens in the input and those in the output: , where and denote the counts of boundary tokens in the input and output, respectively. When the granularity score is equal to 1, the reasoning matches the input in boundary count, suggesting ideal alignment. The consistency score is defined as the similarity between the input sentence and reasoning sentences, i.e., , where and denote the embedding vectors of the reasoning sentences and the input sentence , respectively. We use SentenceBERT (Reimers and Gurevych, 2019) for sequential consistency calculation.333In cases where a single input sentence corresponds to multiple reasoning sentences, we regard them collectively as one segment. We provide the similarity map between the input and the reasoning in the Appendix C.5.
@@ 3.2 Streaming Training Framework
The RoPE (Su et al., 2024) of LLMs represents relative positions by rotating queries and keys, with the attention between a reasoning token and an input token expressed as , where is the input length and are their positional IDs, and is the rotary matrix. However, in streaming scenarios, the concurrent generation of output and reception of input induces positional contention in the encoding space (Guo et al., 2024; Tong et al., 2025). To address this issue, we assign independent position IDs to input and reasoning tokens, both starting from zero. Formally, the positional IDs of reasoning token and input token are both set to , yielding . This design removes positional contention in streaming parallel processing. Furthermore, identical position IDs ensure that, during streaming reasoning, a reasoning sentence is positioned nearest to its associated input sentence and distant from others, which conforms to the essential principle of streaming alignment.
@@ 4.1 Experimental Settings
To evaluate StreamingThinker, we conduct a comprehensive assessment across three representative reasoning tasks: math reasoning, logical reasoning, and context-based QA reasoning. For math reasoning, we selected GSM-symbolic (Mirzadeh et al., 2025) and MetamathQA (Yu et al., 2024). For logical reasoning, we utilized LogicNLI (Tian et al., 2021) and ProofWriter (Tafjord et al., 2020). Finally, for context-based QA reasoning, PubMedQA (Jin et al., 2019) and HotpotQA (Yang et al., 2018) were employed. All datasets were partitioned into dedicated training and testing sets, with detailed specifications provided in the Appendix D.
@@
The evaluation of streaming scenarios can be viewed as a trade-off between performance and latency. For reasoning performance, we adopt Pass@1 score as the accuracy metric to measure the model’s ability to successfully solve problems. Latency is assessed at two levels: token latency and time latency. At the token level, we use token-to-first-token (TTFT) to measure how many input tokens must be observed before the model begins reasoning.444This is similar with time-to-first-token, but measured in terms of token count. For time latency, we set the LLM’s input speed as the average human speaking rate about 150 words per minute (Geva and Yaghoub Zadeh, 2006; Jacewicz et al., 2010), and define the waiting time until the first answer token as the latency.555Detailed definitions and additional evaluation metrics are provided in Appendix F.
@@ 5 Discussion
Prior studies (Fan et al., 2025; Laban et al., 2025) have highlighted the risks of reasoning over incomplete inputs. This phenomenon reflects the inadequacy of batch-processing LLMs when facing local contexts, as their training objectives align with global visibility rather than incremental inference. However, Streaming Thinker circumvents these pitfalls through three distinct mechanisms. First, global information is deferred rather than lost. Unlike scenarios where critical conditions are permanently removed, our framework merely shifts the timing of acquisition, ensuring the model incorporates the full global context after the streaming phase. Second, we employ a conservative reasoning strategy. Instead of attempting premature complex reflection, the model concentrates on shallow reasoning” (e.g., intermediate calculations and entity tracking) during the streaming phase. This functions as an incremental pre-processing step that simplifies raw context. Third, this behavior is enforced via a specialized streaming training paradigm. Unlike batch models that rigidly apply full-context patterns to partial inputs—often falling into the over-thinking” trap—our model is explicitly adapted to local scopes, learning to process available information without jumping to erroneous conclusions.
@@ 6 Related Work
Prior studies on efficient reasoning in LLMs have mainly focused on three directions: token compression (Aggarwal and Welleck, 2025; Xia et al., 2025; Zhang et al., 2025a; Zhao et al., 2026), structural quantization and pruning (Liu et al., 2025; Srivastava et al., 2025; Zhao et al., 2025), and efficient decoding (Pan et al., 2025; Liao et al., 2025). Token compression methods such as Tokenskip (Xia et al., 2025) condense CoT into fewer tokens to reduce the inference cost. Structural approaches leverage quantization or pruning to compress model parameters for efficient reasoning such as (Srivastava et al., 2025). Efficient decoding has been explored through parallel sampling of generation paths (Pan et al., 2025) or by using smaller models for speculative prediction (Liao et al., 2025) to reduce inference latency. In contrast, our work introduces streaming processing as a complementary dimension of efficiency, allowing reasoning to proceed concurrently with input processing to reduce response latency.
@@
Recent research on streaming LLMs has focused on three directions: architecture adaptation (Raffel et al., 2024; Guo et al., 2024; Tong et al., 2025), latency control (Ma et al., 2018; Ahmed et al., 2025; Cheng et al., 2025), and modality adaptation (Chen et al., 2024a; 2025a; Xie and Wu, 2024; Lin et al., 2026). Architecture adaptation addresses the mismatch between decoder-only transformers and streaming settings (Tong et al., 2025), where issues such as positional interference and redundant re-encoding are mitigated through recurrent states, and modified attention mechanisms. Latency control emphasizes fine-grained alignment between input and output, ranging from fixed policies (e.g., wait-k (Ma et al., 2018)) to adaptive scheduling that optimize the accuracy–latency balance. Modality adaptation extends streaming LLMs beyond text to speech recognition (Guo et al., 2024), speech translation (Cheng et al., 2025), and video understanding (Chen et al., 2024a) by integrating modality-specific encoders and synchronization mechanisms. From a reasoning perspective, recent studies (Xie et al., 2025a; Chiang et al., 2025; Xie et al., 2025b) have explored alternating reasoning and generation to approximate streaming operation. Our work differs by explicitly modeling thinking while reading, enabling reasoning to evolve concurrently with incremental input.
@@ B.1 Cognitive Foundations of Streaming Thinking
Our proposed StreamingThinker architecture draws structural inspiration from established models of human discourse comprehension, specifically the Construction-Integration (CI) model proposed by Kintsch (Kintsch, 1988). The CI model posits that comprehension occurs in two alternating cycles:
@@
Inspired by this cognitive process, we proposed in the main text a Streaming Thinking Paradigm that enables large language models to reason concurrently with incremental input. Our streaming phase corresponds to the construction phase, where the model performs ”shallow reasoning” (e.g., entity tracking, intermediate calculation) to process incoming chunks and maintain local coherence (Graesser et al., 1994). Our final reasoning phase corresponds to the integration phase, where the model utilizes the fully accumulated context (KV cache) to synthesize a globally consistent answer. This theoretical alignment explains why our model avoids the ”hallucination” pitfalls of premature guessing: like a human reader, it defers the final integration of complex causal chains until the necessary global information is available.
@@ C.1 Does Incremental Sentence Input Work?
A straightforward idea is to feed sentences into LLMs sequentially and expect sentence-by-sentence reasoning based on the accumulated input. However, our experiments show that when given only a single sentence without streaming fine-tuning, the original reasoning model tends to overthink (Sui et al., 2025). In other words, although this sentence-level data generation method appears to align better with the streaming property, the vanilla model lacks such generative capability and exhibits reduced controllability in CoT generation.
@@ C.4 Quality Evaluation of Streaming CoT
where and denote the embedding vectors of reasoning sentence and input sentence , respectively. We employ SentenceBERT (Reimers and Gurevych, 2019) to obtain embeddings. A higher score reflects that the reasoning faithfully preserves both the order and semantic content of the input stream. In cases where multiple reasoning sentences correspond to one input sentence, we aggregate them as a single segment.